A battery of e.m.f 3.0V is connected across a potentiometer wire AB of 10m.If the jockey J makes contact at the 30cm mark, determine the resistance of length AJ and voltage across AJ

Grade:10

1 Answers

Harshit Singh
3 years ago
Dear student

E = 3.0V;
R / l =10Ω per m
l = 30cm = 0.3m;
R =10Ω per m × 0.3m = 3Ω
Thus , Resistance of AJ = 3 ohms
E / V = l1/ l2
⟹3 / V = 1 / 0.3


=3×0.3
Hence , Voltage across AJ = 0.9V



Thanks